# QE4PE: Word-level Quality Estimation for Human Post-Editing

### v1

@@ Acknowledgements
Gabriele Sarti and Arianna Bisazza acknowledge the support of Imminent, which funded the data collection for this study with a Language Technology Grant. The authors are grateful to Translated S.r.l. and Global Textware B.v. and all translators that collaborated in the data collection process. Gabriele Sarti, Grzegorz Chrupała and Arianna Bisazza also acknowledge the support of the Dutch Research Council (NWO) for the project InDeep (NWA.1292.19.399). Arianna Bisazza is further supported by the NWO Talent Programme (VI.Vidi.221C.009). Ana Guerberof Arenas is supported by the European Research Council (ERC CoG INCREC 101086819).
Broader Impact and Ethical Considerations
Our study explicitly centers the experience of professional translators, responding to recent calls for user-centered evaluation of translation technologies. By prioritizing translators’ perspectives and productivity, we aim to contribute to methods that complement rather than replace human expertise. Our findings highlight a gap between user perception and measured quality improvements, suggesting that future efforts should focus primarily on improving the usability of these methods in editing interfaces. In particular, new assistive approaches for post-editing should not only strive to increase productivity, but rather reduce the cognitive burden associated with post-editing work. This insight is crucial for designing more user-centered quality estimation tools that genuinely support human work. Finally, our results underscores the importance of culturally adaptive design for assessing the effectiveness of proposed methodologies across participants with different cultural backgrounds. All participants in this study were professional translators who provided informed consent. The research protocol ensured anonymity and voluntary participation, with translators recruited through professional translation providers. The study’s open data release further promotes transparency, allowing other researchers to reproduce and extend our findings.
Appendix A Filtering Details and Statistics for QE4PE Data
@@
Documents should contain between 4 and 10 segments, each containing 10-100 words (959 docs). This ensures that all documents are roughly uniform in terms of size and complexity to maintain a steady editing flow (Section 3.5).
@@
The same heuristics were applied to both translation directions, selecting only documents matching our criteria in both cases. Table 3 provide some statistics for the selected examples.
| Task | Domain | # Docs | # Seg. | # Words |
| Pre | Social | 4 | 23 | 539 |
| Biomed. | 2 | 15 | 348 | |
| Main | Social | 30 | 160 | 3375 |
| Biomed. | 21 | 165 | 3384 | |
| Post | Social | 6 | 34 | 841 |
| Biomed. | 2 | 16 | 257 | |
| Total | 64 | 413 | 8744 | |

### v2

@@ Acknowledgements
Authors acknowledge the support of Imminent, which funded the data collection through a Language Technology Grant. Authors are grateful to Translated S.r.l. and Global Textware B.v. and to all translators involved in this study for agreeing to participate in the data collection process. We thank TACL reviewers and editors for their insightful comments. Gabriele Sarti, Grzegorz Chrupała and Arianna Bisazza also acknowledge the support of the Dutch Research Council (NWO) for the project InDeep (NWA.1292.19.399). Arianna Bisazza is supported by the NWO Talent Program (VI.Vidi.221C.009). Ana Guerberof Arenas is supported by the ERC Grant CoG INCREC 101086819.
Appendix A Filtering Details for QE4PE Data
@@
Documents should contain between 4 and 10 segments, each containing 10-100 words (959 docs). This ensures that all documents are roughly uniform in terms of size and complexity to maintain a steady editing flow (Section˜3.5).
@@
The same heuristics were applied to both translation directions, selecting only documents matching our criteria in both cases.